OpenStudy (anonymous):

Can you explain how to factorise 5a+ax-2b+by please

OpenStudy (anonymous):

Put all the lie terms together

OpenStudy (anonymous):

Pout all the like terms together

OpenStudy (anonymous):

a(5+x) + b(-2+y)

OpenStudy (mayankdevnani):

\[\huge 5a+ax-2b+by\] first, seperate this as: \[\huge 5a+ax ~and~-2b+by\] then take common a : \[\huge 5a+ax=a(5+x)\] then take b as common: \[\huge -2b+by=+b(-2+y)\] then combine this: we get, \[\huge a(5+x)+b(-2+y)\] so ^^^^ this is your answer

OpenStudy (mayankdevnani):

got it... @Peanut1

jhonyy9 (jhonyy9):

how to factorise 5a+ax-2b+by so for your help let step by step --- if you check it in the first and second term is common factor a .... ...and how have wrote above @mayankdevnani so is understandably sure right easy

OpenStudy (mayankdevnani):

thank you... @jhonyy9

jhonyy9 (jhonyy9):

yw